A rare but serious disease, D, has been found in 0.01 percent of a certain population. A test has been developed that will be positive, p, for 98 percent of those who have the disease and be positive for 3 percent of those who do not have the disease. Find the probability that a person tested as positive does not have the disease.
Answer #1

Le us denote the events

A :a randomly selected person has the disease

B : a randomly selected person tested as positive

Given

that means P(Ac) = 1 - P(A) = 1 - 0.01 = 0.99

Now,

The probability that a person tested as positive does not have the disease